# Large Scale Masked Autoencoding for Reducing Label Requirements on SAR Data

### v3

@@ 1 Introduction
Satellite remote sensing has fundamentally changed the way we address climate change, offering large scale, high resolution data for applications such as forest mapping Immitzer et al. [2016]Waldeland et al. [2022], wildfire monitoring Hu et al. [2021] and flood detection Tarpanelli et al. [2022]. Data from such tasks is crucial to address problems caused by climate change, but is restricted by the limitations of optical sensing. The inability of these sensors to operate at night, through cloud cover and without atmospheric interference mean that optical data is inappropriate for time sensitive tasks such as natural disaster management Tarpanelli et al. [2022]. Synthetic Aperture Radar (SAR) overcomes these limitations, providing more consistent, all-weather, day-night monitoring capabilities. These enhanced capabilities are invaluable for timely intervention in situations including extreme weather Otto et al. [2016], natural disasters Clarke et al. [2022], rapid ecological shifts Allen et al. [2015], and deforestation Moffette et al. [2021], all of which have implications for climate change.
@@
Self-supervised learning offers the advantage of learning directly from the input data without requiring ground truth labels. Methodologies such as those based on contrastive learning Radford et al. [2021], masked image modelling He et al. [2021] and knowledge distillation Caron et al. [2021] have achieved remarkable successes on RGB image data, yielding significant improvements in various tasks such as image classification, segmentation, and object detection Wang et al. [2022a]Yu et al. [2022], reducing dependency on labelled data. Despite these advancements, the application of self-supervised learning to SAR data remains relatively unexplored Wang et al. [2022b]. Approaches based on data fusion with other remote sensing data sources such as RGB satellite or aerial imagery exist Chen and Bruzzone [2022]Sun et al. [2023], but although these approaches can exploit the specifics of SAR data, they may not be robust to the absence of usable RGB data at night or under cloud cover. A small number of methods operating solely on SAR data exist Ren et al. [2021]Wen et al. [2021]Xu et al. [2021], but have not yet clearly shown the geographic or temporal generalisabilty often lacking in remote sensing models Safonova et al. [2023]. Applying self-supervised learning directly to the large amounts of available unlabelled SAR data would allow practitioners to circumvent the limitations posed by the absence of reliable RGB data at night or in cloudy conditions - improving accuracy and response time in areas such as disaster management and environmental monitoring. Moreover, the use of large-scale, geographically diverse data with a model large enough to accommodate it has the potential to overcome the generalisability issues that often plague remote sensing models, presenting a robust alternative solely based on SAR data.
In this work, we take a self-supervised pretraining scheme - masked autoencoding He et al. [2021] - that has been proven effective on curated RGB imagery, and apply it to polarimetric SAR data on a large set of data covering 8.7% of the Earth’s land surface. We finetune the pretrained model on two downstream tasks - vegetation cover prediction (per-image regression) and land cover classification (semantic segmentation). We show that, in all cases, pretraining improved performance on downstream tasks. We also show that models initialized with pretrained weights still outperform their randomly initialized counterparts when using substantially fewer labels. We show that the pretrained model generalised well to regions that were not seen in the pretraining set.
@@ 2.2 Models
For pretraining we used a masked autoencoder with a ViT-B Dosovitskiy et al. [2021] encoder followed by a reconstruction decoder based on He et al. [2021]. We motivate our model selection by the observation that the original masked autoencoder behaves reasonably with minimal data augmentation He et al. [2021]. Selecting appropriate data augmentations for SAR data introduces additional complexity compared to RGB data - for example, rotation or flipping may introduce invariance to information specific to each polarisation of the instrument. A contrastive approach based on two different SAR modes Radford et al. [2021] - for example, polarimetry with two different polarisations or polarimetry and another mode such as coherence - may similarly neglect information specific to each mode. We therefore choose to omit comparison to additional methods in this short work, although future comparison remains of interest.
@@
For the MODISVEG task, we replaced the reconstruction decoder with a regression head comprising 1D convolutions, of output dimension 196, in both the sequence and hidden dimensions, followed by 3 fully connected layers of sizes {512, 256, 128}. We use ReLU Agarap [2019] activation functions between hidden layers and ELU Clevert et al. [2016] before the regression output.
For the ESAWC task, we followed SETR-PUP Zheng et al. [2021]. We increase the number of decoder layers compared to the original implementation to maintain a maximum upsampling of 2×2$\times$2 × per layer.
@@ A.1 Data Split
[…]
@@ A.2 SAR Reconstructions
Reconstructions by the masked autoencoder of SAR data masked during pretraining can be seen in Figure A.2. Note that the explicit aim of pretraining is to learn input features, not to obtain high reconstruction accuracy. The model largely predicts low-frequency features, as in He et al. [2021].
@@ A.3 Training Details
Hyperparameter details for MAE-based pretraining and task finetuning can be seen in Table A.2. The patch size was halved relative to the size of the image compare to He et al. [2021], based on the intuition that distant pixels in remote sensing imagery are less likely to be correlated than distant pixels in curated photographs. This intuition appeared to be evidenced by linear probe performance on the validation data for vegetation prediction, although we do not report the results from the linear probe as it was not used on all tasks or regions. Beyond this, the probe was not used to make further design decisions. Increasing or decreasing the learning rate by an order of magnitude did not improve convergence on the validation data. No significant hyperparameter tuning was undertaken beyond these two decisions, as the computational expense of performing an equal level of tuning for all tasks and regions was too high.
@@
Encoder (Params) ViT-B Dosovitskiy et al. [2021] (88.8M) ViT-B Dosovitskiy et al. [2021] (88.8M) ViT-B Dosovitskiy et al. [2021] (88.8M)
Decoder (Params) Reconstruction He et al. [2021] (5.5M) $1\times 1$ Conv + FC (Section 2.2; 19.7M) SETR-PUP Zheng et al. [2021] (3.0M)
@@ A.4 Training Histories
Aauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v4

@@ 1 Introduction
Satellite remote sensing has fundamentally changed the way we address climate change, offering large scale, high resolution data for applications such as forest mapping [1][2], wildfire monitoring [3] and flood detection [4]. Data from such tasks is crucial to address problems caused by climate change, but is restricted by the limitations of optical sensing. The inability of these sensors to operate at night, through cloud cover and without atmospheric interference mean that optical data is inappropriate for time sensitive tasks such as natural disaster management [4]. Synthetic Aperture Radar (SAR) overcomes these limitations, providing more consistent, all-weather, day-night monitoring capabilities. These enhanced capabilities are invaluable for timely intervention in situations including extreme weather [5], natural disasters [6], rapid ecological shifts [7], and deforestation [8], all of which have implications for climate change.
@@
Self-supervised learning offers the advantage of learning directly from the input data without requiring ground truth labels. Methodologies such as those based on contrastive learning [9], masked image modelling [10] and knowledge distillation [11] have achieved remarkable successes on RGB image data, yielding significant improvements in various tasks such as image classification, segmentation, and object detection [12][13], reducing dependency on labelled data. Despite these advancements, the application of self-supervised learning to SAR data remains relatively unexplored [14]. Approaches based on data fusion with other remote sensing data sources such as RGB satellite or aerial imagery exist [15][16], but although these approaches can exploit the specifics of SAR data, they may not be robust to the absence of usable RGB data at night or under cloud cover. A small number of methods operating solely on SAR data exist [17][18][19], but have not yet clearly shown the geographic or temporal generalisabilty often lacking in remote sensing models [20]. Applying self-supervised learning directly to the large amounts of available unlabelled SAR data would allow practitioners to circumvent the limitations posed by the absence of reliable RGB data at night or in cloudy conditions - improving accuracy and response time in areas such as disaster management and environmental monitoring. Moreover, the use of large-scale, geographically diverse data with a model large enough to accommodate it has the potential to overcome the generalisability issues that often plague remote sensing models, presenting a robust alternative solely based on SAR data.
In this work, we take a self-supervised pretraining scheme - masked autoencoding [10] - that has been proven effective on curated RGB imagery, and apply it to polarimetric SAR data on a large set of data covering 8.7% of the Earth’s land surface. We finetune the pretrained model on two downstream tasks - vegetation cover prediction (per-image regression) and land cover classification (semantic segmentation). We show that, in all cases, pretraining improved performance on downstream tasks. We also show that models initialized with pretrained weights still outperform their randomly initialized counterparts when using substantially fewer labels. We show that the pretrained model generalised well to regions that were not seen in the pretraining set.
@@ 2.2 Models
For pretraining we used a masked autoencoder with a ViT-B [21] encoder followed by a reconstruction decoder based on [10]. We motivate our model selection by the observation that the original masked autoencoder behaves reasonably with minimal data augmentation [10]. Selecting appropriate data augmentations for SAR data introduces additional complexity compared to RGB data - for example, rotation or flipping may introduce invariance to information specific to each polarisation of the instrument. A contrastive approach based on two different SAR modes [9] - for example, polarimetry with two different polarisations or polarimetry and another mode such as coherence - may similarly neglect information specific to each mode. We therefore choose to omit comparison to additional methods in this short work, although future comparison remains of interest.
@@
For the MODISVEG task, we replaced the reconstruction decoder with a regression head comprising 1D convolutions, of output dimension 196, in both the sequence and hidden dimensions, followed by 3 fully connected layers of sizes {512, 256, 128}. We use ReLU [22] activation functions between hidden layers and ELU [23] before the regression output.
For the ESAWC task, we followed SETR-PUP [24]. We increase the number of decoder layers compared to the original implementation to maintain a maximum upsampling of 2×2$\times$2 × per layer.
@@ A.1 Data Split
We used repeated geographic bands to define our training, validation and test sets. These bands can be seen for the four AOIs in Figure A.1. Coverage was determined by intersection with the coverage of the ARIA S1 GUNW dataset [^4] https://asf.alaska.edu/data-sets/derived-data-sets/sentinel-1-interferograms/, which was not used in this work. This approach minimises data leakage compared with a fully randomised split, while also reducing the train-test distribution shift that would occur when using one geographically contiguous band for each set. Data was split into the training, validation and test sets at a 60:20:20 ratio. A total of 737,050 tiles were generated, spanning an area of 1.4793×$\times 10^{7}$km2. See Table A.1 for a breakdown by AOI.
@@ A.2 SAR Reconstructions
Reconstructions by the masked autoencoder of SAR data masked during pretraining can be seen in Figure A.2. Note that the explicit aim of pretraining is to learn input features, not to obtain high reconstruction accuracy. The model largely predicts low-frequency features, as in [10].
@@ A.3 Training Details
Hyperparameter details for MAE-based pretraining and task finetuning can be seen in Table A.2. The patch size was halved relative to the size of the image compare to [10], based on the intuition that distant pixels in remote sensing imagery are less likely to be correlated than distant pixels in curated photographs. This intuition appeared to be evidenced by linear probe performance on the validation data for vegetation prediction, although we do not report the results from the linear probe as it was not used on all tasks or regions. Beyond this, the probe was not used to make further design decisions. Increasing or decreasing the learning rate by an order of magnitude did not improve convergence on the validation data. No significant hyperparameter tuning was undertaken beyond these two decisions, as the computational expense of performing an equal level of tuning for all tasks and regions was too high.
@@
Encoder (Params) ViT-B [21] (88.8M) ViT-B [21] (88.8M) ViT-B [21] (88.8M)
Decoder (Params) Reconstruction [10] (5.5M) $1\times 1$ Conv + FC (Section 2.2; 19.7M) SETR-PUP [24] (3.0M)